# Generalization Bounds via Meta-Learned Model Representations: PAC-Bayes and Sample Compression Hypernetworks

### v2

@@ 4.2 Numerical Results on a Synthetic Problem
Figure 3: Illustration of decision boundaries generated by downstream predictors $h_{\gamma}$, when parameters $\gamma$ are given by a PAC-Bayes decoder $\operatorname{\mathscr{D}}_{\psi}({\boldsymbol{\mu}})$ with input ${\boldsymbol{\mu}}=(\mu_{1},\mu_{2})\in{\mathbb{R}}^{2}$. Left: The first latent dimension is fixed ($\mu_{1}=0.97$), while the second varies ($\mu_{2}\in[-1,1]$). Right: the first dimension varies ($\mu_{1}\in[-1,1]$), while the second is fixed $\mu_{2}=-0.27$. The encoder generated the message (0.97, -0.27).
@@
4.3 Test Case: Noisy MNIST
Table 1: Comparison of different meta-learning methods on the MNIST-pixels-swap task. The 95% confidence interval is reported for generalization bound and test error, computed over 20 test tasks. The best result in each column is bolded.
@@ 4.3 Test Case: Noisy MNIST
We compare our approaches to algorithms yielding PAC-Bayesian bounds as benchmarks. We chose a fully connected network with no more than 3 hidden layers and a linear output layer as a backbone, as per the selected benchmarks. We also report the performances of a strawman: an opaque encoder outputting nothing, followed by a decoder which input is a predefined sole constant scalar; though the reconstructor can be trained, the hypernetwork always generates the same predictor, no matter the input. We report the test bounds and test errors on the novel tasks of various methods in Table 1. More details about the experiment setup and the hyperparameters can be found in Appendix H.
When it comes to generalization bounds, our approaches outperform the benchmarks. However, our approaches cannot learn from the task environment, their performances being similar to those of the strawman. This is because the difference between the various tasks is essentially noise; such a setting is not well-suited for our approaches, which rely on the encapsulation of the differences between tasks in a message or a compression set. The benchmark methods perform well because the posterior (over all of the downstream predictor), for each test task, is similar to the prior. Indeed, the various tasks do not vary much; this is reflected by a single predictor (the strawman, an opaque hypernetwork) yielding competitive performances across all of the test tasks.
@@ 4.4 Test Case: Binary MNIST and CIFAR100 Tasks
Table 2: Comparison of different meta-learning methods on the MNIST (CIFAR100) binary task. The 95% confidence interval is reported for generalization bound and test error, computed over 34 (50) test tasks. The best result in each column is bolded.
@@
In light of the analysis made in the previous subsection, we now explore an experimental setup where a prior model cannot encompass most information of the various tasks. To do so, we create a variety of binary tasks involving the various classes of the MNIST (CIFAR100) dataset, where a task corresponds to a random class versus another one. We create 90 (150) such tasks, where the meta-test set corresponds to all of the tasks involving either label 0 or label 1, chosen at random, leading to a total of 34 (50) meta-test tasks. Each training task contains 2000 (1200) examples from the train split of the MNIST (CIFAR100) original task, while each test task contains at most 2000 (200) examples from the test split of the original task. We consider the same benchmarks as previously. We report the test bounds and test errors on the novel tasks of various methods in Table 2 with their corresponding latent representation information in Table 3.
As expected, in such a setting, when it comes to the benchmarks, the posterior for each task is required to be truly different from the prior in order to perform well (as attested by the strawman’s test error, being similar to a random guess). Thus, the many benchmarks generate uninformative generalization bounds, even though their test loss is competitive. On the other hand, many of our approaches achieve competitive empirical performances while also having informative generalization bounds, since the downstream predictors can be truly different from one another without impacting the quality of the bound.
The architectures reported in Table 3, along with the empirical performances in Table 2, confirm that 1) the encoder / the message compressor and the sample compressor are able to correctly distill the particularities of the task at hand via a message or a compression set, and 2) that the reconstructor is able to utilize this representation to judiciously generate the downstream predictor.
See Table 4 in Appendix H for results involving the bound from Theorem 2.3 and Theorem 2.5.
@@ 5 Conclusion
We developed a new paradigm for deriving generalization bounds in meta-learning by leveraging the PAC-Bayesian framework or the Sample Compression theory. We also present a new generalization bound that permits the coupling of both paradigms. We develop meta-learning hypernetworks based on these results. We show that PAC-Bayes approaches of the literature do not scale when the various tasks in a meta-learning environment have important discrepancies while our approaches still yield low losses and tight generalization bounds.
The approaches we presented could be enhanced by having the model dynamically select the compression set size and the latent representation (or message) size. Finally, since the bound values are not impacted by the complexity of the decoder or the downstream predictor, our approach could be used to finetune predictors as complex as large language models (LLMs) while still providing tight generalization bounds.
@@
$~{}~{}~{}~{}\mathcal{L}_{D}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))\leq\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-C\widehat{\mathcal{L}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))+\frac{\ln(P_{J}(\operatorname{\mathbf{j}})\cdot P_{M(|\mathbf{j}|)}(\omega)\cdot\delta)}{m-\operatorname{|\mathbf{j}|}})]$
@@
$\forall\mathbf{j}\in J,\omega\in M(|\mathbf{j}|)~{}:~{}\mathcal{L}_{D}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))\leq\widehat{\mathcal{L}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))+\frac{\lambda\varsigma^{2}}{2}-\frac{\ln(P_{J}(\operatorname{\mathbf{j}})\cdot P_{M(|\mathbf{j}|)}(\omega)\cdot\delta)}{\lambda(m-|\operatorname{\mathbf{j}}|)}$
@@ Appendix B General PAC-Bayes Sample Compression Theorems
We derive two new PAC-Bayes Sample compression theorems for real-valued losses, both for the loss on the validation set and on the train set. Theorem B.1 extends the setting of Laviolette & Marchand (2005, 2007); Thiemann et al. (2017), whilst Theorem B.2 extends the setting of Germain et al. (2011, 2015).
@@ B.1 First General PAC-Bayes Sample Compression Theorem
We first present a new general PAC-Bayes Sample compression theorem, from which we will derive a sample compression bound for continuous messages. Interestingly, this theorem is directly at the intersection between PAC-Bayes theory and sample compression theory. Indeed, if we restrict the compressions to have no compression set, this bound reduces to Theorem 2.1. Moreover, if we restrict the model to a discrete family $\Omega$ and to Dirac measures as posteriors on $J$ and $\Omega$, this bound is almost exactly reduced to Theorem 2.3, albeit being possibly slightly less tight, as the complexity term denominator $m-\operatorname{|\mathbf{j}|}$ of the latter is replaced by the worst case: $m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}$
@@
$\Delta\quantity(\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))\leq\frac{1}{m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}}\quantity[\mathrm{KL}(Q||P)+\ln\quantity(\frac{\mathcal{A}_{\Delta}(m)}{\delta})]$
@@
$\leq\mathrm{KL}(Q||P)+\ln\quantity(\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))})$ (Change of measure)
@@
$\eta\Delta\quantity(\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))\leq\mathrm{KL}(Q||P)+\ln\quantity(\frac{1}{\delta}\operatorname*{\mathbb{E}}_{T\sim\operatorname{\mathcal{D}}^{m}}\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))})$
@@
$=\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}$ (i.i.d. assumption)
We can then set
${\mathcal{A}}_{\Delta}(m)=\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}$ (10)
Finally, for all $\operatorname{\mathbf{j}}\in J$, we need to bound the moment generating function
@@
To bound the moment generating function using the usual PAC-Bayes techniques, we need $\eta\leq m-\operatorname{|\mathbf{j}|}$ for all $\operatorname{\mathbf{j}}\in J$. Thus, the largest value of $\eta$ (which gives the tightest bound) that can be used is $\eta=m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}$. ∎

### v3

@@ 4.2 Numerical Results on a Synthetic Problem
Table 1: Comparison of different meta-learning methods on the MNIST-pixels-swap task. The 95% confidence interval is reported for generalization bound and test error, computed over 20 test tasks. The best (smallest) result in each column is bolded.
@@ 4.3 Test Case: Noisy MNIST
We compare our approaches to algorithms yielding PAC-Bayesian bounds as benchmarks. The reported bounds concern the generalization property of the trained model on a given test task. We chose a fully connected network with no more than 3 hidden layers and a linear output layer as a backbone, as per the selected benchmarks. We also report the performances of a strawman: an opaque encoder outputting nothing, followed by a decoder which input is a predefined sole constant scalar; though the reconstructor can be trained, the hypernetwork always generates the same predictor, no matter the input. We report the test bounds and test errors on the novel tasks of various methods in Table 1. More details about the experiment setup can be found in Appendix H.
When it comes to generalization bounds, our approaches outperform the benchmarks. However, our approaches cannot learn from the task environment, their performances being similar to those of the strawman. This is because the difference between the various tasks is very subtle; such a setting is not well-suited for our approaches, which rely on the encapsulation of the differences between tasks in a message or a compression set. The benchmark methods perform well because the posterior (over all of the downstream predictor), for each test task, is similar to the prior. Indeed, the various tasks do not vary much; this is reflected by a single predictor (the strawman, an opaque hypernetwork) yielding competitive performances across all of the test tasks. This matter is empirically confirmed by the next experiment.
@@ 4.4 Test Case: Binary MNIST and CIFAR100 Tasks
Table 2: Comparison of different meta-learning methods on the MNIST and CIFAR100 binary tasks. The 95% confidence interval is reported for generalization bound and test error, computed over test tasks. The best (smallest) result in each column is bolded.
@@
Figure 7: Test error and generalization bound for the PB SCH algorithm as a function of both the compression set size and the message size on binary MNIST tasks. The 95% confidence interval is reported for both the generalization bound and the test error, computed over 34 test tasks.
In light of the analysis made in the previous subsection, we now explore an experimental setup where a prior model cannot encompass most of the information of the various tasks. To do so, we create a variety of binary tasks involving the various classes of the MNIST (CIFAR100) dataset, where a task corresponds to a random class versus another one. We create 90 (150) such tasks, where the meta-test set corresponds to all of the tasks involving either label 0 or label 1, chosen at random, leading to a total of 34 (50) meta-test tasks. Each training task contains 2000 (1200) examples from the train split of the MNIST (CIFAR100) original task, while each test task contains at most 2000 (200) examples from the test split of the original task. We consider the same benchmarks as previously. We report the test bounds and test errors on the novel tasks of various methods in Table 2 with their corresponding latent representation information in Table 3.
Table 4: KL value of two benchmarks on the pixel swap tasks and the binary MNIST task. The 95% confidence interval is reported for generalization bound and test error, computed over 20 test tasks.
Algorithm 100 Pixels swap 200 Pixels swap 300 Pixels swap binary MNIST
(Pentina & Lampert, 2014) 3.833 $\pm$ 0.444 5.760 $\pm$ 0.720 9.604 $\pm$ 1.035 14.02 $\pm$ 0.018
(Zakerinia et al., 2024) 63.91 $\pm$ 24.12 159.9 $\pm$ 54.78 223.9 $\pm$ 58.60 661.2 $\pm$ 20.32
As expected, in such a setting, when it comes to the benchmarks, the posterior for each task is required to be truly different from the prior in order to perform well (as attested by the strawman’s test error, being similar to a random guess). We present in Table 4 the penalty (KL) value measuring the distance between the prior and posterior for two benchmarks on both the pixel swap experiments and the binary MNIST variant. There is a significant gap between the KL value reported for the various pixel swap tasks and the binary MNIST task, which confirms our insight. Thus, the benchmarks methods generate uninformative generalization bounds, even though their test loss is competitive. On the other hand, most of our approaches achieve competitive empirical performances while also having informative generalization bounds, since the downstream predictors can be truly different from one another without impacting the quality of the bound.
The architectures reported in Table 3, along with the empirical performances in Table 2, confirm that 1) the encoder / the message compressor and the sample compressor correctly distill the particularities of the task at hand, and 2) that the reconstructor is able to utilize this representation to judiciously generate the downstream predictor.
Figure 7 depicts the test error and generalization bound for our PB SCH algorithm as a function of both the compression set size and the message size. We recall that Tables 1 and 2 report the performances of the models obtaining the best validation error. Figure 7 helps to grasp the inner workings of our proposed approach: using a larger message seems better-suited for minimizing the test error, but a trade-off between the compression set size and the message size is required to obtain the best bounds. Interestingly, when the message size $|{\boldsymbol{\mu}}|$ is restricted to be small, we clearly see the benefit of using a compression set ($c>0$).
See Table 5 in Appendix H for results involving the bound from Theorem 2.3 and Theorem 2.5. See also in Appendix H a decomposition of the various terms involved in the composition of our bounds and the benchmarks’.
@@ 5 Conclusion
We developed a new paradigm for deriving generalization bounds in meta-learning by leveraging the PAC-Bayesian framework or the Sample Compression theory. We also present a new generalization bound that permits the coupling of both paradigms. We develop meta-learning hypernetworks based on these results. We show that many PAC-Bayes approaches do not scale when the various tasks in an environment have important discrepancies while our approaches still yield low losses and tight generalization bounds.
The approaches we presented could be enhanced by having the model dynamically select the compression set size and the latent representation (or message) size or using a larger architecture for the reconstruction function, inspired from MotherNet (Mueller et al., 2024). Finally, since the bound values are not impacted by the complexity of the decoder or the downstream predictor, our approach could be used to get tight generalization bounds for very large models of the multi-billion parameter scale, assuming that the parameters varying across tasks admit a compact representation (e.g., a LoRA adapter).
@@
Acknowledgements
This research is supported by the NSERC/Intact Financial Corporation Industrial Research Chair in Machine Learning for Insurance. Pascal Germain is supported by the Canada CIFAR AI Chair Program, and the NSERC Discovery grant RGPIN-2020-07223. Mathieu Bazinet is supported by a FRQNT B2X scholarship (343192, doi: https://doi.org/10.69777/343192).
@@ Acknowledgements
In the setting of Theorem 2.3, with
$\Delta_{C}(q,p)=-\ln(1-(1-e^{-C})p)-Cq\quad\mbox{(where $C>0$)\,,}$
with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{m}$, we have
@@
$~{}~{}~{}~{}\mathcal{L}_{D}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))\leq\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-C\widehat{\mathcal{L}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))+\frac{\ln(P_{J}(\operatorname{\mathbf{j}})\cdot P_{M(|\mathbf{j}|)}(\omega)\cdot\delta)}{m-\operatorname{|\mathbf{j}|}})].$
@@
In the setting of Theorem 2.3, with
$\Delta(q,p)=\mathrm{kl}(q,p)=q\cdot\ln\frac{q}{p}+(1-q)\cdot\ln\frac{1-q}{1-p}\,,$
with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{m}$, we have
@@
$\forall\mathbf{j}\in J,\omega\in M(|\mathbf{j}|)~{}:~{}\mathcal{L}_{D}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))\leq\widehat{\mathcal{L}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))+\frac{\lambda\varsigma^{2}}{2}-\frac{\ln(P_{J}(\operatorname{\mathbf{j}})\cdot P_{M(|\mathbf{j}|)}(\omega)\cdot\delta)}{\lambda(m-|\operatorname{\mathbf{j}}|)}.$
@@ Appendix B General PAC-Bayes Sample Compression Theorems
We derive two new PAC-Bayes Sample compression theorems for real-valued losses, both for the loss on the complement set and on the train set. Theorem B.1 extends the setting of Laviolette & Marchand (2005); Thiemann et al. (2017), whilst Theorem B.2 extends the setting of Germain et al. (2011, 2015).
@@ B.1 First General PAC-Bayes Sample Compression Theorem
We first present a new general PAC-Bayes Sample Compression theorem, from which we will derive a sample compression bound for continuous messages. Interestingly, this theorem is directly at the intersection between PAC-Bayes theory and sample compression theory. Indeed, if we restrict the model to have no compression set, this bound reduces to Theorem 2.1. Moreover, if we restrict the model to a discrete family $\Omega$ and to Dirac measures as posteriors on $J$ and $\Omega$, this bound is almost exactly reduced to Theorem 2.3, albeit being possibly slightly less tight, as the complexity term denominator $m-\operatorname{|\mathbf{j}|}$ of the latter is replaced by the worst case: $m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}$
@@
$\Delta\quantity(\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))\leq\frac{1}{m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}}\quantity[\mathrm{KL}(Q||P)+\ln\quantity(\frac{\mathcal{A}_{\Delta}(m)}{\delta})],$
@@
$\leq\mathrm{KL}(Q||P)+\ln\quantity(\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))}).$ (Change of measure)
@@
$\eta\Delta\quantity(\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim Q}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))\leq\mathrm{KL}(Q||P)+\ln\quantity(\frac{1}{\delta}\operatorname*{\mathbb{E}}_{T\sim\operatorname{\mathcal{D}}^{m}}\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\!\!\!\!\!\!e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}).$
@@
$=\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}.$ (i.i.d. assumption)
For all $\operatorname{\mathbf{j}}\in J$, we need to bound the moment generating function
@@
To bound the moment generating function using the usual PAC-Bayes techniques, we need $\eta\leq m-\operatorname{|\mathbf{j}|}$ for all $\operatorname{\mathbf{j}}\in J$. Thus, the largest value of $\eta$ (which gives the tightest bound) that can be used is $\eta=m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}$.
As the exponential is monotonically increasing and $m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}\leq m-\operatorname{|\mathbf{j}|}\forall\operatorname{\mathbf{j}}$, we have
$\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{\eta\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}\leq\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{T_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}e^{(m-\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{j}}},\omega)))}\eqqcolon{\mathcal{A}}_{\Delta}(m).$